# Strict quantization for compact pseudo-Kähler manifolds and group actions

### v1

@@
Abstract
The asymptotic results for Berezin-Toeplitz operators yield a strict quantization for the algebra of smooth functions on a given Hodge manifold. It seems natural to generalize this picture for quantizable pseudo-Kähler manifolds in presence of a group action. Thus, in this setting we introduce a Berezin transform which has a complete asymptotic expansion on the preimage of the zero set of the moment map. It leads in a natural way to prove that certain quantization maps are strict.
Keywords: CR manifolds, Toeplitz operators, group actions, star products
Mathematics Subject Classification: 32Vxx, 32A25, 53D50
@@ 1 Introduction
This paper is a continuation of [GH1], here we investigate further properties of the star product induced by Toeplitz operators, let us briefly recall some known results. In [BMS], the authors studied Berezin Toeplitz quantization for Hodge manifolds. In [MM2], see also references therein, Toeplitz operators for symplectic manifolds are introduced. In [GH1] we considered Berezin-Toeplitz quantization for a class of CR manifolds and investigated the properties of the corresponding star product, see also [GH2] for further results in this geometric setting. In [S2], additional properties for the Berezin Toeplitz star product are proved: it is shown that the properties of Berezin Toeplitz operators lead in a natural way to prove that certain quantization maps are strict. In this paper we aim to generalize these results to pseudo-Kähler manifolds in the presence of a group action.
Berezin-Toeplitz quantization has been studied extensively, giving rise to a vast literature. Without claiming to be exhaustive, let us mention a few references most relevant to the present discussion: [BoGu], [BPU], [C], [G], [HM], [KS], [P], [MM1], [MM2], [Z]. In particular, we refer to the surveys [S2], [W] for considerations about the induced star product. More recently, in [HHMS] and [Sh], the authors study Toeplitz operators associated with classical first-order pseudodifferential operators and apply them to problems in CR geometry. Eventually, in connection with the principle that “quantization commutes with reduction”, we refer to [Ma].
We now provide a more precise formulation of the concepts introduced above. Let $(X,\,T^{1,0}X)$ be a compact, connected and orientable CR manifold of dimension $12n+$ 1, $n\geq 1$, where $T^{1,0}X$ is a subbundle of the complexified tangent bundle $TX\otimes\mathbb{C}$. There is a unique sub-bundle $HX$ of $TX$ such that
$HX\otimes\mathbb{C}=T^{1,0}X\oplus T^{0,1}X$
which is called the horizontal bundle. Let $J\,:\,HX\rightarrow HX$ be the complex structure map given by
$J(u+\overline{u})=iu-i\overline{u}\,,$
for every $u\in T^{1,0}X$. By complex linear extension of $J$ to $TX\otimes\mathbb{C}$, the $i$-eigenspace of $J$ is $T^{1,0}X$.
Since $X$ is orientable, there always exists a real non-vanishing $1$-form
$\omega_{0}\in{C}^{\infty}(X,T^{*}X)$
so that $\langle\,\omega_{0}(x)\,,\,u\,\rangle=0$, for every $u\in H_{x}X$, for every $x\in X$. The one form $\omega_{0}$ is called contact form and it naturally defines a volume form $\mathrm{dV}_{X}$ on $X$. For each $x\in X$, we define a quadratic form on $HX$ with respect to $\omega_{0}$ by
${L}_{x}(U,V)=\frac{1}{2}\mathrm{d}\omega_{0}(JU,V),\qquad\forall\ U,V\in H_{x}X.$
Then, we extend ${L}$ to $HX\otimes\mathbb{C}$ by complex linear extension; for $U,V\in T^{1,0}_{x}X$,
${L}_{x}(U,\overline{V})=\frac{1}{2}\,\mathrm{d}\omega_{0}(JU,\overline{V})=-\frac{1}{2i}\,\mathrm{d}\omega_{0}(U,\overline{V}).$
The Hermitian quadratic form ${L}_{x}$ on $T^{1,0}_{x}X$ is called Levi form at $x$ (with respect to $\omega_{0}$). Hereafter we shall always assume that the Levi form is non-degenerate with $n_{-}$ negative and $n_{+}$ positive eigenvalues, $n_{+}+n_{-}=n$. The Reeb vector field $R\in{C}^{\infty}(X,TX)$ is defined to be the non-vanishing vector field determined by
$\omega_{0}(R)\equiv-1,\quad\mathrm{d}\omega_{0}(R,\cdot)\equiv 0\ \ \mbox{on $TX$}.$
There is a Reeb vector field such that the flow of it induces a transversal CR $\mathbb{R}$-action on $X$. Suppose that all orbits of the flow of $R$ are compact. Then $X$ admits a transversal CR circle action, which we denote by $e^{i\,\theta}\cdot$, where $\theta\in\mathbb{R}$.
Remark 1.1.
Note that the quotient, denoted by $M$, of $X$ by the circle action is a pseudo-Kähler manifold, see for example [BoGa]. Indeed, $M$ is endowed with an integral symplectic form $\omega$ such that
$2\,\pi^{*}\omega=\mathrm{d}\omega_{0}$
where $\pi\,:\,X\rightarrow M$ is the projection. The horizontal bundle $HX$ is the pullback of the tangent bundle $TM$ via $\pi$ and $J$ induces a complex structure $J_{M}$ on $M$. Thus,
$g:=\omega(\,\cdot,\,J_{M}\,\cdot\,)$
is easily seen to be a pseudo-Riemannian structure on $M$.
Vice versa, suppose that $(M,\omega,J_{M})$ is a compact pseudo-Kähler manifold with $[\omega]$ defining an integral cohomology class. Then, there exists a line bundle $(L,h)$ on $M$ such that the unique holomorphic connection compatible with $h$ has curvature $-i\,\omega$. Then, the circle bundle $X$ inside the dual of the line one is a contact manifold satisfying the assumption of this article.
To be consistent with [GH1] and [GH2] we prefer to work on $X$ and make explicit reference to $M$ later on.
Assume that $X$ admits an action of a compact connected Lie group $G$ of dimension $d$. In this work, we assume that the $G$ action preserves $\omega_{0}$ and $J$. That is, $g^{\ast}\omega_{0}=\omega_{0}$ on $X$ and $g_{\ast}J=Jg_{\ast}$ on $HX$, for every $g\in G$, where $g^{*}$ and $g_{*}$ denote the pull-back map and push-forward map of $G$, respectively.
Let $\mathfrak{g}$ denote the Lie algebra of $G$. For any $\xi\in\mathfrak{g}$, we write $\xi_{X}$ to denote the vector field on $X$ induced by $\xi$. That is,
$(\xi_{X}u)(x)=\frac{\partial}{\partial t}\left[u\left(e^{t\xi}\,x\right)\right]_{|_{t=0}},$
for any $u\in{C}^{\infty}(X)$.
@@
Now, we define the quantizing space. Fix $g\in G$ and $x\in X$, for any $r\in\mathbb{N}$ let
$g^{*}:\Lambda^{r}_{x}(T^{*}X\otimes\mathbb{C})\rightarrow\Lambda^{r}_{g^{-1}\circ x}(T^{*}X\otimes\mathbb{C})$
be the pull-back map. Since $G$ preserves $J$, we have
$g^{*}:T^{*0,q}_{x}X\rightarrow T^{*0,q}_{g^{-1}\circ x}X\,$
for all $x\in X$, where the bundle $T^{*0,q}X$ is the $q$-th exterior power of $(T^{(0,1)}X)^{*}$, whose space of sections is denoted by $\Omega^{0,q}(X)$ and its elements are called $(0,q)$-forms. Thus, for $u\in\Omega^{0,q}(X)$, we have $g^{*}u\in\Omega^{0,q}(X)$. Put
$\Omega^{0,q}(X)^{G}:=\left\{u\in\Omega^{0,q}(X);\,g^{*}u=u,\ \ \forall g\in G\right\}\,.$
@@
$L^{2}_{(0,q)}(X)^{G}:=\left\{u\in L^{2}_{(0,q)}(X):\,g^{*}u=u,\ \ \forall g\in G\right\}.$
Put
$({\rm Ker\,}\Box^{q}_{b})^{G}:={\rm Ker\,}\Box^{q}_{b}\cap L^{2}_{(0,q)}(X)^{G}.$
The $G$-invariant Szegő projection is the orthogonal projection
$S^{(q)}_{G}:L^{2}_{(0,q)}(X)\rightarrow({\rm Ker\,}\Box^{q}_{b})^{G}$
@@
$S^{(q)}_{G}(x,y)\in{D}^{\prime}(X\times X,T^{*0,q}X\boxtimes(T^{*0,q}X)^{*})$
be the distributional kernel of $S^{(q)}_{G}$, ${D}^{\prime}$ denotes the space of distribution sections, see Section 2.2 for notation.
In this paper, we will assume that $\Box^{q}_{b}$ has $L^{2}$ closed range. For simplicity we assume that $n_{+}\neq n_{-}$ and we fix $q=n_{-}$. Let $f\in C^{\infty}(X,\,\mathbb{C})$ be invariant under the circle action and the action of $G$, then the $G$-invariant Toeplitz operator is given by
$T^{G}[f]:=S^{(n_{-})}_{G}\circ f\cdot\circ S^{(n_{-})}_{G}:L^{2}_{(0,n_{-})}(X)\rightarrow({\rm Ker\,}\Box^{n_{-}}_{b})^{G}\,.$
The transversal circle action can be used to define the corresponding Fourier components. First, we assume that
$e^{i\theta}\cdot(g\,x)=g\,(e^{i\theta}\cdot x),\ \ \mbox{for every $x\in X$, $\theta\in[0,2\pi[$, $g\in G$}.$
Let $u\in\Omega^{0,q}(X)$ be arbitrary. Define
$Ru:=\frac{\partial}{\partial\theta}\bigr{(}(e^{i\theta})^{*}u\bigr{)}|_{\theta=0}\in\Omega^{0,q}(X).$

### v2

@@
[…]
Strict Quantization for Compact Pseudo-Kähler Manifolds and Group Actions
[…]
Strict Quantization for Compact Pseudo-Kähler
Manifolds and Group Actions
[…]
Andrea GALASSO
[…]
A. Galasso
[…]
Dipartimento di Matematica e Applicazioni, Università degli Studi di Milano Bicocca,
[…]
Received January 29, 2025, in final form June 17, 2025; Published online June 25, 2025
[…]
The asymptotic results for Berezin–Toeplitz operators yield a strict quantization for the algebra of smooth functions on a given Hodge manifold. It seems natural to generalize this picture for quantizable pseudo-Kähler manifolds in presence of a group action. Thus, in this setting we introduce a Berezin transform which has a complete asymptotic expansion on the preimage of the zero set of the moment map. It leads in a natural way to prove that certain quantization maps are strict.
[…]
CR manifolds; Toeplitz operators; star products; group actions
[…]
32V20; 32A25; 53D50
@@ 1 Introduction
This paper is a continuation of [7]. Here, we further investigate the properties of the star product induced by Toeplitz operators, with a focus on establishing conditions under which certain quantization maps are strict. The present paper aims to extend known results for Kähler manifolds to the setting of pseudo-Kähler manifolds equipped with a group action. We emphasize that the purpose of this introduction is to provide a broad overview of the results.
Consider a compact connected complex manifold $(M,J)$ of real dimension $2n$, where $J$ is an integrable complex structure. Let $\omega$ be a real, non-degenerate closed $2$-form on $M$, and assume that it satisfies the compatibility condition $\omega(JX,JY)=\omega(X,Y)$, $\forall X,Y\in TM$. In this case, the triple $(M,\omega,J)$ is called a pseudo-Kähler manifold. Unlike in the standard Kähler case, we do not assume that $\omega$ is positive definite. Instead, $\omega$ has constant signature $(n_{-},n_{+})$, meaning that at each point $m\in M$, the symmetric bilinear form $g_{m}(X,Y):=\omega_{m}(X,JY)$ has signature ${(n_{-},n_{+})}$.
Let $(L,h_{L})$ be a Hermitian line bundle over $M$, endowed with a Hermitian connection $\nabla^{L}$ whose curvature $R^{L}$ satisfies $R^{L}=-2\pi{\rm i}\omega$.
Definition 1.1.
For each positive integer $k$, the quantization space $\mathcal{Q}_{k}(M)$ of $(M,\omega)$ with respect to the polarization $J$ is defined as the space of harmonic $(0,n_{-})$-forms with values in the $k$-th tensor power $L^{\otimes k}$ of $L$,
$\mathcal{Q}_{k}(M):=\ker\square^{(n_{-})}_{k},$
where $\square^{(n_{-})}_{k}$ is the Kodaira Laplacian acting on $L^{\otimes k}$-valued $(0,n_{-})$-forms.
In the special case when $n_{-}=0$ (i.e., when $g$ is positive definite), this construction recovers the classical geometric quantization: $\mathcal{Q}_{k}(M)$ coincides with the space $H^{0}\bigl{(}M,L^{\otimes k}\bigr{)}$ of holomorphic sections. Hence, this framework generalizes the Kähler case to pseudo-Kähler manifolds with indefinite metrics. There is a natural inner product on $\mathcal{Q}_{k}(M)$, we refer to Section 2 for the definition and for more details.
@@
Definition 1.2.
A star product for the algebra $C^{\infty}(M)$ is given by the formal power series
$g\star h=\sum_{j=0}^{+\infty}C_{j}(g,h)\nu^{-j}$
such that $\star$ is an associative $\mathbb{C}[[\nu]]$-linear product, that is, $(g\star h)\star l=g\star(h\star l),$ for all $g,h,l\in C^{\infty}(M)$ and $C_{0}(g,h)=g\cdot h$, $C_{1}(g,h)-C_{1}(h,g)={\rm i}\{g,h\}$, for all $g,h\in C^{\infty}(M)$. The star product is said to be of Wick type if and only if the function appearing as first argument is differentiated in holomorphic directions while the one appearing as second argument is differentiated in anti-holomorphic directions.
In addition to the defining conditions, a star product may enjoy further properties, such as
  • (1)
    (Unit) Let $1$ be the constant function equals to $1$ everywhere, then $1\star f=f\star 1=f$ for every smooth function on $M$.
  • (2)
    (Parity) For every smooth functions $f$ and $g$ on $M$, then $\overline{f\star g}=\overline{f}\star\overline{g}$.
  • (3)
    (Trace) Denote the trace on $\mathcal{Q}_{k}(M)$ by $\mathrm{Tr}_{k}$. It induces a $\mathbb{C}[[\nu]]$-linear map
    $\mathrm{Tr}\colon\ {C}^{\infty}(M)[[\nu]]\rightarrow\nu^{-n}\mathbb{C}[[\nu]]$
    such that $\mathrm{Tr}(f\star g)=\mathrm{Tr}(g\star f)$.
Let $G$ be a compact connected Lie group having dimension $\dim_{\mathbb{R}}G=n_{G}$. Assume that $M$ admits a Hamiltonian and holomorphic action of $G$ with moment map $\Phi$. Suppose that 00, in the dual of the Lie algebra, is a regular value for $\Phi$. Assume that the pseudo-metric $g$ restricted to the orbit through any $m\in\Phi^{-1}(0)$ is non-degenerate and the action of $G$ on $\Phi^{-1}(0)$ is free. Then the symplectic reduction $M_{G}:=\Phi^{-1}(0)/G$ is a quantizable pseudo-Kähler manifold, see, for example, [31, Appendix B]. More precisely, let $\iota\colon\Phi^{-1}(0)\hookrightarrow M$ denote the inclusion map. Then the following holds:
  • (1)
    the topological quotient $M_{G}:=\Phi^{-1}(0)/G$ is a smooth manifold of dimension $2n-2n_{G}$, and the quotient map $\pi\colon\Phi^{-1}(0)\to M_{G}$ is a principal $G$-bundle;
  • (2)
    there exist a unique pseudo-Riemannian metric $g_{G}$ and complex structure $J_{G}$ on $M_{G}$ such that $\pi^{*}g_{G}=\iota^{*}g$, $\pi^{*}J_{G}=\iota^{*}J$, and the 2-form $\omega_{G}:=g_{G}(\cdot,J_{G}\cdot)$ defines a symplectic form on the quotient.
Furthermore, we assume that the symmetric bilinear form $g_{G}$ has signature $\bigl{(}n_{-}^{G},n_{+}^{G}\bigr{)}$.
Suppose that the action of $G$ on $M$ lifts on $(L,h_{L})$. Then there exists an unitary representation of $G$ on $\mathcal{Q}_{k}(M)$. Let $\mathcal{Q}_{k}(M)^{G}$ be the space of $G$-fixed vectors in $\mathcal{Q}_{k}(M)$. In Section 3, for every smooth function $f$ on $M_{G}$, we define a $G$-invariant Toeplitz operators $T^{G}_{k}[f]$ acting on $\mathcal{Q}_{k}(M)^{G}$. Our next result show that $G$-invariant Toeplitz operators induces a star product $\star$ which defines a strict deformation quantization in the sense of [18, Chapter 2]. The following theorem is a consequence of Theorem 2.3 which is formulated further down.
Theorem 1.3.
There exists a unique formal star-product of Wick type
$f\star g:=\sum_{j=0}^{+\infty}\nu^{j}C_{j}(f,g),\qquad C_{j}(f,g)\ \text{lies in}\ C^{\infty}(M_{G}),$
such that for every smooth functions $f$ and $g$ on $M_{G}$, for every $N\in\mathbb{N}$ we have with suitable constants $K_{N}(f,g)$, and for all $k\in\mathbb{N}$
$\left\lVert T^{G}_{k}[f]T^{G}_{k}[g]-\sum_{j=0}^{N-1}k^{-j}T^{G}_{k}[C_{j}(f,g)]\right\rVert=K_{N}(f,g)k^{-N}.$ (1.1)
Furthermore, the star product $\star$ satisfies the properties $(1)$, $(2)$, and $(3)$ of Definition 1.2.
To a certain extent the strategy to show Theorem 1.3 is similar to the strategy done in [32], details are given in Section 7. Now, if we put $G=\{e\}$, where $e$ is the identity element, we recover previous results in [32], and we generalize it to pseudo-Kähler manifolds. Furthermore, we refer to the surveys [15, 33], and [38] for detailed discussions on the induced star products.
In order to better contextualize the aim of this work and to motivate our main results, we recall some relevant developments in the literature. In [1], the authors studied Berezin–Toeplitz quantization on Hodge manifolds, their proofs are based on the theory of generalized Toeplitz structures developed by L. Boutet de Monvel, V. Guillemin, and J. Sjöstrand in the framework of microlocal analysis, see [2]. In [33], additional properties of the Berezin–Toeplitz star product were established, and it was shown that the structure of Berezin–Toeplitz operators naturally leads to the strictness of certain quantization maps. The idea of considering pseudo-Kähler manifolds is suggested by results contained in [13, 23], in particular see [22, Theorem 1.7] and [23, Section 8.2], where the authors worked out the asymptotic expansion for the Bergman kernel for $(0,q)$-forms under the assumption that the curvature is non-degenerate.
In the geometric quantization of pseudo-Kähler manifolds, two main strategies are available: one based on the theory of Berezin–Toeplitz quantization with vector bundles, as developed by X. Ma and G. Marinescu [23], and another one employing Toeplitz operators on the associated circle bundle. In this work, we adopt the second approach. This choice is primarily motivated by the fact that it aligns with the framework we have developed in a previous article [7].
The main approach of X. Ma and G. Marinescu is the local index theory, especially the analytic localization techniques developed by J.-M. Bismut and G. Lebeau. Our approach is based of Fourier integral operators of complex type as developed by A. Melin and J. Sjöstrand. It is worth noting that, in principle, both methods could be employed in the context of this article. Our method can be applied to study Berezin–Toeplitz star product for Reeb invariant smooth functions on non-degenerate CR manifolds with transversal and CR $\mathbb{R}$-action.
We now introduce the setting for our quantization commutes with reduction result in the pseudo-Kähler case. Suppose that the action of $G$ on $M$ lifts to an action on $(L,h_{L})$. Then $(L,h_{L})$ descends to a Hermitian line bundle $\bigl{(}L_{G},h_{L}^{G}\bigr{)}$ over $M_{G}$ with curvature $R^{L_{G}}=-2\pi{\rm i}\omega_{G}$. Recall that the standard Kähler case is recovered when $n_{-}=0$, and consequently $n_{-}^{G}=0$.
This allows us to define the quantization of the reduced space as follows.
Definition 1.4.
The quantization space $\mathcal{Q}_{k}(M_{G})$ is defined as the space of harmonic $\bigl{(}0,n_{-}^{G}\bigr{)}$-forms with values in $L_{G}^{\otimes k}$
$\mathcal{Q}_{k}(M_{G}):=\ker\Box_{k}^{(n_{-}^{G})},$
where $\square^{(n_{-}^{G})}_{k}$ is the Kodaira Laplacian acting on $L^{\otimes k}_{G}$-valued $\bigl{(}0,n_{-}^{G}\bigr{)}$-forms.
There is a canonical map $\mathcal{Q}_{k}(M)^{G}\to\mathcal{Q}_{k}(M_{G})$, and our main result is the following (the proof is given in Section 5.2).
Theorem 1.5 (quantization commutes with reduction).
Under the above assumptions, for every $k\in\mathbb{N}$, the canonical map $\mathcal{Q}_{k}(M)^{G}\longrightarrow\mathcal{Q}_{k}(M_{G})$ is an isomorphism of finite-dimensional Hilbert spaces.
The principle that “quantization commutes with reduction” was originally formulated by V. Guillemin and S. Sternberg in the Kähler setting [9]. The present result extends this principle to the pseudo-Kähler setting, of which the Kähler case is a particular instance.
In the context of general symplectic manifolds, proofs of this principle were developed by E. Meinrenken [26], M. Vergne [36, 37], and E. Meinrenken and R. Sjamaar [27]. Independently, Y. Tian and W. Zhang [35] gave a different analytic approach. For a survey on the topic, we refer to [21], where the connection with the CR setting introduced in [12] is also highlighted.
Finally, let us also mention that in many of the aforementioned approaches, the quantization space is defined as the index of a twisted Spinc-Dirac operator. In contrast, our proof of Theorem 1.5 follows the geometric strategy introduced in [9], adapted to the setting of indefinite curvature. We refer to Section 5.1 for the detailed construction and background.
2 Background and statement of the results
2.1 Examples
There are many examples of pseudo-Kähler manifolds that serve as natural candidates for quantization. First, we refer to the thesis [31], where the author investigates the symplectic and pseudo-Riemannian geometry of the $\mathbb{P}{\rm SL}(3,\mathbb{R})$ Hitchin component associated with a closed orientable surface. When the genus is at least $2$, the author defines a mapping class group-invariant pseudo-Kähler metric on the Hitchin component.
A second example is given by the Grauert tube of an analytic pseudo-Riemannian manifold; see [39] for an introduction in the classical Riemannian setting. Finally, we present a natural example arising in the context of the Heisenberg group.
Before describing this example, we recall some material from [6]. A CR manifold $\bigl{(}X,T^{1,0}X\bigr{)}$, with contact form $\omega_{0}$, is said to be nondegenerate if its Levi form is nondegenerate at every point, for precise definitions, we refer the reader to Section 2.3. In this case, replacing $\omega_{0}$ by $-\omega_{0}$ if necessary, we may assume that the (constant) dimension $k$ of a maximal positive-definite subspace for the Levi form is at least $n/2$. We then say that $X$ is $k$-strongly pseudoconvex. When $k=n$, we simply say that $X$ is strongly pseudoconvex. If $X$ is $k$-strongly pseudoconvex with $k<n$, then by [6, Lemma 13.1], there exist smooth subbundles $E^{+}$ and $E^{-}$ of $T^{1,0}X$ such that $T^{1,0}X=E^{+}\oplus E^{-}$, the Levi form is positive definite on $E^{+}$, negative definite on $E^{-}$, and ${E^{+}\perp E^{-}}$ with respect to the Levi form.
Example 2.1.
Let $\mathbb{H}_{n}=\mathbb{C}^{n}\times\mathbb{R}$ denote the Heisenberg group. Occasionally, as in [5], it is useful to modify the group law on $\mathbb{H}_{n}^{{\rm pol}}=\mathbb{R}^{2n+1}$ via $(p,q,t)\cdot(p^{\prime},q^{\prime},t^{\prime})=(p+p^{\prime},q+q^{\prime},t+t^{\prime}+pq^{\prime})$ where ${z=p+{\rm i}q}$. Moreover, following [6, pp. 436–437], we may define a left-invariant CR structure on $\mathbb{H}_{n}^{{\rm pol}}$ which is $k$-strongly pseudoconvex, with $n/2\leq k\leq n$, by taking $T^{1,0}\mathbb{H}_{n}^{{\rm pol}}$ to be the bundle spanned by $Z_{1},\dots,Z_{k},\overline{Z}_{k+1},\dots,\overline{Z}_{n}$, where $Z_{j}=\frac{\partial}{\partial z_{j}}+{\rm i}\overline{z}_{j}\frac{\partial}{\partial t}$.
By compactifying this example as in [5, p. 68], one obtains a natural compact pseudo-Kähler manifold. Let $\Gamma$ be the subgroup of $\mathbb{H}_{n}^{{\rm pol}}$ consisting of points with integer coordinates
$\Gamma=\big{\{}(p,q,t)\in\mathbb{H}^{\rm pol}_{n}\mid p,q\in\mathbb{Z}^{n}\text{ and }t\in\mathbb{Z}\big{\}}.$
Then $\Gamma$ is a discrete subgroup of $\mathbb{H}_{n}^{{\rm pol}}$, and the right coset space $M=\Gamma\backslash\mathbb{H}_{n}^{{\rm pol}}$ is a compact manifold. It is easy to verify that the half-open unit cube $Q^{2n+1}$ in $\mathbb{H}_{n}^{{\rm pol}}$ is a fundamental domain for $\Gamma$, meaning that each right coset of $\Gamma$ contains precisely one point of $Q^{2n+1}$. Hence, topologically, $M$ can be viewed as the closed unit cube $Q^{2n+1}$ with certain boundary faces identified—a kind of “twisted torus”. Moreover, by the above discussion, $M$ inherits a pseudo-Kähler structure.
2.2 Toeplitz operators in the pseudo-Kähler setting
@@
where $T^{*0,q}M$ is the bundle of $(0,q)$ forms. Consider the vector bundle $T^{*0,q}M\otimes L^{\otimes k}$ whose space of smooth sections is denoted by $\Omega^{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}$.
[…]
$(s_{1},s_{2})_{k}=\int_{M}\langle s_{1}|s_{2}\rangle_{h^{L^{\otimes k}}}\mathrm{dV}_{M},\qquad s_{1},s_{2}\in\Omega^{0,q}\bigl{(}M,L^{\otimes k}\bigr{)},$
[…]
$\square^{(q)}_{k}\colon\ \Omega^{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}\rightarrow\Omega^{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}$
be the Kodaira Laplacian. We denote by the same symbol the $L^{2}$ extension of $\square^{(q)}_{k}$
$\square^{(q)}_{k}\colon\ \operatorname{Dom}\square^{(q)}_{k}\subset L^{2}_{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}\rightarrow L^{2}_{0,q}\bigl{(}M,L^{\otimes k}\bigr{)},$
where $\operatorname{Dom}\square^{(q)}_{k}=\big{\{}u\in L^{2}_{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}\mid\square^{(q)}_{k}u\in L^{2}_{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}\big{\}}$. The projection
$P^{(q)}_{k}\colon\ L^{2}_{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}\rightarrow H^{q}\bigl{(}M,L^{\otimes k}\bigr{)}:=\operatorname{Ker}\square^{(q)}_{k}$
onto the kernel of $\square^{(q)}_{k}$ is called Bergman projector or Bergman projection. Since $\square^{(q)}_{k}$ is elliptic, its distributional kernel satisfies
$P^{(q)}_{k}(\cdot,\cdot)\in\mathcal{C}^{\infty}\bigl{(}M\times M,L^{k}\otimes\bigl{(}\bigl{(}T^{*0,q}M\bigr{)}\boxtimes\bigl{(}T^{*0,q}M\bigr{)}^{*}\bigr{)}\otimes\bigl{(}L^{k}\bigr{)}^{*}\bigr{)}.$
Now, we shall follow [14] and define Toeplitz operators in this framework. Fix a smooth function $f$ on $M$. Define the Berezin–Toeplitz operator
$T^{(q)}_{k}[f]:=P^{(q)}_{k}\circ f\circ P^{(q)}_{k}\colon\ L^{2}_{0,q}\bigl{(}M,L^{\otimes k}\bigr{)}\rightarrow H^{q}\bigl{(}M,L^{\otimes k}\bigr{)}:=\operatorname{Ker}\square^{(q)}_{k};$
we always fix $q=n_{-}$ and we write $T_{k}[f]=T^{(n_{-})}_{k}[f]$. Note that if $M$ is compact complex manifold endowed with a positive line bundle $L$, we put $q=n_{-}=0$ and we recover the standard Kähler setting. Thus, the quantizing Hilbert space is $\operatorname{Ker}\square^{(q)}_{k}$, for Riemann–Roch-type results we refer to [14] and [8, Section 1.2].
2.3 CR formulation
The circle bundle $X$ inside the dual of the line bundle $L$ is a compact, connected and orientable CR manifold of dimension $12n+$ 1. We now provide a more precise formulation of these concepts. Let $\bigl{(}X,T^{1,0}X\bigr{)}$ be a compact, connected and orientable CR manifold of dimension $12n+$ 1, $n\geq 1$, where $T^{1,0}X$ is a subbundle of the complexified tangent bundle $TX\otimes\mathbb{C}$. There is a unique sub-bundle $HX$ of $TX$ such that $HX\otimes\mathbb{C}=T^{1,0}X\oplus T^{0,1}X$ which is called the horizontal bundle. Let ${J\colon HX\rightarrow HX}$ be the complex structure map given by $J(u+\overline{u})={\rm i}u-{\rm i}\overline{u}$, for every ${u\in T^{1,0}X}$. By complex linear extension of $J$ to $TX\otimes\mathbb{C}$, the ${\rm i}$-eigenspace of $J$ is $T^{1,0}X$.
Since $X$ is orientable, there always exists a real non-vanishing $1$-form $\omega_{0}\in{C}^{\infty}(X,T^{*}X)$ so that $\langle\omega_{0}(x),u\rangle=0$, for every $u\in H_{x}X$, for every $x\in X$. The one form $\omega_{0}$ is called contact form and it naturally defines a volume form $\mathrm{dV}_{X}$ on $X$. For each $x\in X$, we define a quadratic form on $HX$ with respect to $\omega_{0}$ by ${L}_{x}(U,V)=\frac{1}{2}{\rm d}\omega_{0}(JU,V)$, $\forall U,V\in H_{x}X$. Then, we extend ${L}$ to ${HX\otimes\mathbb{C}}$ by complex linear extension; for $U,V\in T^{1,0}_{x}X$,
${L}_{x}\bigl{(}U,\overline{V}\bigr{)}=\frac{1}{2}{\rm d}\omega_{0}\bigl{(}JU,\overline{V}\bigr{)}=-\frac{1}{2{\rm i}}{\rm d}\omega_{0}\bigl{(}U,\overline{V}\bigr{)}.$
The Hermitian quadratic form ${L}_{x}$ on $T^{1,0}_{x}X$ is called Levi form at $x$ (with respect to $\omega_{0}$). Hereafter we shall always assume that the Levi form is non-degenerate with $n_{-}$ negative and $n_{+}$ positive eigenvalues, $n_{+}+n_{-}=n$. The Reeb vector field $R\in{C}^{\infty}(X,TX)$ is defined to be the non-vanishing vector field determined by $\omega_{0}(R)\equiv-1$, ${\rm d}\omega_{0}(R,\cdot)\equiv 0$ on $TX$. There is a Reeb vector field such that the flow of it induces a transversal CR $\mathbb{R}$-action on $X$. Suppose that all orbits of the flow of $R$ are compact. Then $X$ admits a transversal CR circle action, which we denote by ${\rm e}^{{\rm i}\theta}\cdot$, where $\theta\in\mathbb{R}$. Note that this is the case of the unit circle bundle ${X\subset L^{\vee}}$, with projection $\pi\colon X\rightarrow M$ and connection contact form $\omega_{0}$ such that ${{\rm d}\omega_{0}=2\pi^{*}(\omega)}$. Then $(X,\omega)$ is a CR manifold and $\omega_{0}$ is the contact form.
[…]
$S^{(q)}(x,y)\in{D}^{\prime}\bigl{(}X\times X,T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$
be the distributional kernel of $S^{(q)}$, ${D}^{\prime}$ denotes the space of distribution sections, see Section 3.2 for notation. In this paper, we will assume that $\Box^{q}_{b}$ has $L^{2}$ closed range. This hypothesis is always satisfied for the circle bundle case we are considering.
Fix $q=n_{-}$ and let $f\in C^{\infty}(X,\mathbb{C})$, then the Toeplitz operator is given by
$T^{X}[f]:=S^{(n_{-})}\circ f\cdot\circ S^{(n_{-})}\colon\ L^{2}_{(0,n_{-})}(X)\rightarrow\operatorname{Ker}\Box^{n_{-}}_{b}.$
The transversal circle action can be used to define the corresponding Fourier components which we denote by $T_{k}^{X}[f]$, see [7] where we use pseudodifferential operators of order zero (furthermore we refer to [34] for composition with certain classical pseudodifferential operators of order one). Let $u\in\Omega^{0,q}(X)$ be arbitrary. Define
$Ru:=\frac{\partial}{\partial\theta}\bigl{(}\bigl{(}{\rm e}^{{\rm i}\theta}\bigr{)}^{*}u\bigr{)}|_{\theta=0}\in\Omega^{0,q}(X).$